OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 18 raised to the power of 1 over 5 is called the fifth root of 18. (5 points) Part B: The length of a rectangle is 2 units and its width is square root of 2 unit. Is the area of the rectangle rational or irrational? Justify

OpenStudy (anonymous):

i dont understand how to do any of this what is it even asking me??? ????

OpenStudy (anonymous):

I'm not really sure but 18^1 = 18. And typically the roots of something are in the denominator, so (18^1)/5 is 18/5 and 5 would be the 5th of 18 possible roots?

OpenStudy (anonymous):

Not sure but that's my guess

OpenStudy (anonymous):

Part A: Use the properties of exponents to explain why 18 raised to the power of 1 over 5 is called the fifth root of 18. this is by the rational exponents property. a¹/ⁿ = ⁿ√a Part B: The length of a rectangle is 2 units and its width is square root of 2 unit. Is the area of the rectangle rational or irrational? Justify your answer. => irrational L = 2 => rational, W = √2 => irrational area of rectangle = L*W = rational*irrational =